Eighteen years ago Alan was twice as old as Tony. Alan is eight years older than Tony. Find Tony's age now.

Answer by lwsshak3(11628)

Eighteen years ago Alan was twice as old as Tony. Alan is eight years older than Tony. Find Tony's age now.
**
let x=tony's age now
x+8=alan's age now
..
18 years ago
x-18=tony's age
x+8-18=x-10
..
x-10=2(x-18)
x-10=2x-36
x=26
tony's age now=26